# A Comprehensive Survey of Small Language Models in the Era of Large Language Models: Techniques, Enhancements, Applications, Collaboration with LLMs, and Trustworthiness

### v1

@@ 3.2. Supervised Fine-Tuning (SFT) for Enhancing SLM performance
Supervised Fine-Tuning (SFT) employs a training methodology similar to pre-training but is specifically tailored to align models to adhere to the instructions encapsulated within various instructional datasets. This approach is designed to refine the model’s responsiveness and appropriateness to given contexts as dictated by the training data. For example, various models, such as Alpaca (Taori et al., 2023), UltraChat (Ding et al., 2023), WizardLM (Xu et al., 2023b), SlimOrca (Lian et al., 2023), ShareGPT (Wang et al., 2024b), Capybara (Daniele and Suphavadeeprasit, 2023), Deita (Liu et al., 2023b), and MetaMathQA (Yu et al., 2024a), incorporates a suite of conversational datasets to enhance their capabilities in context-aware dialogue and instruction adherence. Usually, as shown in Figure 9, existing SFL methods can be categorized into 3 categories: (i) Classical fine-tuning with downstream data (Devlin et al., 2019; Radford et al., 2019) trains SLMs on task-specific annotated data, transferring general language representations to specific tasks such as sentiment analysis. In the LLM era, this approach remains effective, such as enhancing LLMs by calibrating responses or assigning risk scores with smaller models such as BERT (Zhao et al., 2023b), or optimizing for mobile devices with MobileBERT (Sun et al., 2020). (ii) Instruction tuning with LLM-generated data (Taori et al., 2023; Ding et al., 2023; Lian et al., 2023) or human-generated questions with LLM annotations (Wang et al., 2024b) aims to align generative models with specific instructions, enhancing their instruction-following and reasoning capabilities. For example, Alpaca 7B (Taori et al., 2023) uses 52k ChatGPT-generated instruction-following examples from 175 self-instructed seed tasks to tune Llama 7B (Touvron et al., 2023a). Meanwhile, StableLM (Bellagente et al., 2024; Tow et al., [n. d.]) is trained on the Restruct-v1 dataset, which includes summarization, question-answering, and sentiment analysis tasks, using instruction data from (Longpre et al., 2023b). (iii) Preference optimization with human feedback (Rafailov et al., 2024; Wang et al., 2024b; Ouyang et al., 2022) aims to better align language models with human preferences. Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) gathers human-preferred data, trains a reward model, and fine-tunes the LM using reinforcement learning. Direct Preference Optimization (DPO) (Rafailov et al., 2024) provides a simpler alternative to RLHF. Unlike RLHF, DPO avoids explicit reward modeling and reinforcement learning techniques. Instead, it adjusts the log probabilities of preferred versus non-preferred responses using a dynamic weighting mechanism, preventing model degradation issues typical of methods relying on probability ratios. For instance, Llama 3.2 1B & 3B apply SFT and DPO in post-training to enhance alignment with instructions and human preferences.
@@ 3.3. Data Quality in Knowledge Distillation (KD)
Augment Data from Other Models. Due to the limitations of model size, studies have shown that training SLMs requires simple and comprehensible data (Eldan and Li, 2023; Xi et al., 2024; Lee et al., 2024). TinyStory (Eldan and Li, 2023) demonstrates that language models with a relatively small number of parameters (tens of millions) can still generate coherent stories tailored for children aged 3-4 years. This is achieved by prompting GPT-3.5 or GPT-4 (Achiam et al., 2023) to create simple and easily understandable stories from three keywords selected from a foundational vocabulary of 1,500 words. The generated stories are then used to train SLMs, enabling them to produce similar narratives. This approach shows that varied and comprehensible data can help smaller models exhibit behaviors similar to those of larger language models, such as obeying scaling laws and achieving enhanced performance. Many efforts to enhance the Chain-of-Thought (CoT) capabilities of small models involve using LLMs to generate high-quality CoT data. These data are then employed to train small models in an end-to-end fashion to mimic the CoT reasoning process (Xi et al., 2024; Ma et al., 2023b). AS-ES Learning (Xi et al., 2024) argues that previous methods often overlook the limited capacity of small models to learn complex reasoning, despite being provided with very detailed reasoning processes. Even these detailed processes still require more nuanced capabilities, such as extraction and abstraction. Therefore, this study introduces a novel training paradigm that categorizes reasoning steps into extractive segments, which remind the model of the context and set the stage for subsequent conclusions, and abstractive segments that infer additional insights not explicitly stated in the context.
@@
Figure 11. Distillation Techniques for Enhancing SLM Performance
@@ 3.5. Performance Improvement through Quantization
Quantization-Aware Training (QAT) differs from PTQ in that it includes a training phase after the model has been quantized. When models are quantized to extremes, such as 2-bit or 1-bit, performance typically drops significantly, but further training can help the model retain its capabilities. For instance, to mitigate performance degradation from binarization, PB-LLM (Shang et al., 2023) selectively binarizes only non-salient weights, preserving the most salient ones at higher precision. This method effectively reduces the model size without significantly impacting performance. Salient weights are chosen based on their magnitude, ensuring that the most influential weights maintain higher precision to preserve the model’s reasoning capabilities. The paper explores both post-training quantization (PTQ) and quantization-aware training (QAT) to fine-tune and recover the performance of partially binarized models, achieving a balance between compression and accuracy. OneBit (Xu et al., 2024) and BitNet (Wang et al., 2023e) address the severe performance degradation associated with 1-bit quantization by decomposing floating-point matrices and employing mixed-precision strategies. Specifically, OneBit introduces Sign-Value-Independent Decomposition (SVID), which decomposes a floating-point matrix into a 1-bit matrix and two floating-point vectors. This method allows LLMs to be quantized to a 1-bit level while minimizing performance loss. By retaining critical information with the floating-point vectors, OneBit effectively balances extreme compression with maintaining model accuracy. BitNet b1.58 (Ma et al., 2024) improves on the original BitNet by introducing a ternary matrix weight system -1, 0, 1, resulting in a 1.58-bit model. BitNet b1.58 matches the performance of full-precision models starting from a 3 billion parameter size while further reducing memory and latency costs. LLM-QAT (Liu et al., 2023a) employs data-free knowledge distillation, where the pre-trained model itself generates data for fine-tuning the quantized model (student) using logit distillation from the full-precision model (teacher). This method incorporates quantization of weights, activations, and key-value cache, achieving accurate 4-bit quantization for weights and key-value caches, and 6-bit for activations, demonstrating substantial improvements over existing post-training quantization methods.
@@ 3.6. Techniques in LLMs Contributing to SLMs
To enhance the performance of LLMs, various techniques such as Retrieval-Augmented Generation (RAG) and Mixture of Experts (MoE) are employed. This section discusses their potential to maintain or improve the performance of SLMs within constrained computational budgets. However, effectively integrating these advanced techniques into SLMs, which have inherently limited capabilities, remains an open challenge.
Retrieval Augmented Generation (RAG) enhances the capabilities of language models in knowledge-intensive tasks by incorporating a retrieval mechanism. This approach allows models to access relevant contextual information from a data repository in response to user queries. By integrating this retrieved data, RAG-equipped models gain a better understanding of specific topics, enabling more informed and accurate outputs. For SLMs, a significant concern is whether they possess the capacity for long-context reasoning. A recent study (Yazan et al., 2024) compares the performance of the original FP16 and the quantized INT4 on multiple 7B and 8B LLMs, indicating that for a 7B LLM already performing effectively, quantization does not compromise its performance or its ability to reason over long contexts. Another study (Liu et al., 2024g) compares SLMs at the 7B level with RAG to larger models such as GPT-3.5 and GPT-4 (Achiam et al., 2023), suggesting that SLMs equipped with Retrieval-Augmented Generation can sometimes perform comparably or even better than LLMs. These findings indicate that RAG for SLMs is effective and represents a promising direction for future research.
Mixture-of-Experts (MoE) offers an efficient scaling strategy for LLMs through expert parallelism. However, this approach requires significant memory overhead, necessitating model compression techniques for its adoption in SLMs. Several studies (Kim et al., 2023a; Li et al., 2024d) have found that traditional quantization techniques are less effective for MoE models. The Mixture of Quantized Experts (MoQE(Kim et al., 2023a) demonstrates that expert layers in MoE models are more resilient to quantization compared to traditional feedforward network (FFN) layers. They apply quantization only to expert weights and achieve superior performance compared to dense models trained on the same dataset. Similarly, another study (Li et al., 2024d) suggests that since FFN weights only engage with a subset of input tokens, more bits should be allocated to attention weights. These studies collectively indicate that in MoE architectures, FFN layers should be either quantized or allocated fewer bits.
Insights: Here are some insights from LLM techniques that contribute to SLMs: Implementing RAG in SLMs shows significant promise. Developing techniques to better tailor retrieved information for SLMs could be particularly beneficial. Quantization methods can effectively compress large MoE models by reducing bit allocation on FFN layers. Future research might explore the potential of training small MoE models from scratch.
@@ 4. Applications of Small Language Models
In this section, we delve into the applications of small language models (SLMs) across various NLP tasks and their deployment strategies. Due to benefits such as enhanced privacy, faster inference, and lower memory requirements, many NLP applications are now leveraging SLMs over LLMs, employing specialized techniques to enhance SLM performance. Additionally, deploying SLMs often involves considerations of memory and runtime efficiency, which are crucial for optimizing resource use on budget-constrained edge devices, particularly mobile phones. Then, we will discuss task-specific applications of SLMs and their deployment methods on mobile and edge devices.
@@ 4.1. Task-specific SLM Applications
SLM in QA Alpaca (Taori et al., 2023) Tune Llama 7B (Touvron et al., 2023a) using 52k ChatGPT-generated examples from 175 seed tasks.
Stable Beluga 7B (Mahan et al., [n. d.]) Employ explanation tuning to Llama-2 7B (Touvron et al., 2023b) on an Orca-style dataset with explanatory LLM answers. capabilities
Fine-tuned BioGPT Guo et al. (2023a) Fine-tuning BioGPT (1.6B) (Luo et al., 2022) on PubMedQA.
Financial SLMs (Phogat et al., 2024) Transfer financial knowledge from GPT-4 (Achiam et al., 2023) to multiple SLMs, such as Phi-3-Mini (Abdin et al., 2024), via program of thought (PoT).
ColBERT (Gichamba et al., 2024) Fetch retrieval documents for SLMs to answer complex domain-specific questions.
@@
Rationale Ranking (Hartill et al., 2023a) For OOD questions, generate intermediate reasoning steps or rationales.
SLM in Coding Phi-3.5-mini (Abdin et al., 2024) New addition to the Phi-3 series and focus on high-quality, reasoning-dense data.
TinyLlama (Zhang et al., 2024i) A transformer model with 1.1 billion parameters, has been trained from scratch on a massive corpus of 3 trillion tokens.
CodeLlama (Roziere et al., 2023) A derivative of Llama 2, undergo a rigorous fine-tuning process on domain-specific datasets
CodeGemma (Team, 2024a) Stemming from Google DeepMind’s Gemma framework, also exhibit a focused approach to enhancing coding capabilities through fine-tuning.
@@
SLIM (Wang et al., 2024f) Step-by-step Knowledge Distillation
@@
SLM in Web Search Content encoder (Chang et al., [n. d.]; Humeau et al., [n. d.]; Lu et al., 2020) Encode concatenated queries and documents and integrates correlations at character, word, and phrase levels post-output layer.
Ranker (Chu et al., 2022; Nogueira and Cho, 2019) Introduce a retrieval framework that identifies candidates, which are subsequently re-ranked using a specially fine-tuned T5
Rewriter (Ma et al., 2023d) Utilize the ”rewrite-retrieve-read” framework to bridge the gap between queries and needed knowledge by rewriting inputs.
@@
Revolutionizing Mobile Interaction (Carreira et al., 2023) Text-to-action control and tests on 6GB Android devices and 4GB devices
@@
On-device Agent for Text Rewriting (Zhu et al., 2023b) Data Knowledge Distillation from LLMs
@@ 4.1.1. SLM Applications in Question-Answering
Question-answering (QA) is a fundamental task in the NLP field, demanding language models to exhibit abilities in understanding language, reasoning, common sense, and recalling specialized knowledge. Typically, larger language models yield better QA performance. However, the substantial size of these models introduces challenges such as immense computational requirements, privacy concerns when using proprietary LLMs, and difficulties in customization. These issues lead researchers and developers to favor small language models (SLMs) in scenarios that demand efficiency, privacy, and customization. Therefore, we explore methods to enhance the capabilities of SLMs in QA across three key areas: (i) Instruction Tuning of Generic SLMs for QA, (ii) Instruction Tuning of Domain-Specific SLMs for QA, and (iii) Enhancing SLMs for Out-of-Domain Questions.
@@
Instruction Tuning Generic SLMs for QA. Despite the Phi series’ impressive question-answering capabilities, the cost of training with over 3.4T tokens on 512 H100 GPUs for 10 days (Abdin et al., 2024) poses a challenge for many researchers and developers. Instruction tuning (Wei et al., 2022) offers a cost-effective alternative fine-tuning method, enhancing small models via fine-tuning on large foundation models’ outputs. Alpaca 7B (Taori et al., 2023) tunes Llama 7B (Touvron et al., 2023a) using 52k ChatGPT-generated examples from 175 seed tasks. In blind tests, the performance of Alpaca 7B matches the LLM text-davinci-003 (Brown et al., 2020), winning 90 of 179 comparisons. This behavior cloning is effective in mimicking the style of teacher models. However, this approach may not improve small model performance for reasoning-intensive QA tasks where accuracy, not style, is crucial (Chia et al., 2024). To counter this, Stable Beluga 7B (Mahan et al., [n. d.]) employs explanation tuning, refining Llama-2 7B (Touvron et al., 2023b) on an Orca-style dataset with explanatory LLM answers to enhance reasoning capabilities. Explanation Tuning extracts detailed answers from LLMs using system instructions, but effectiveness varies with system instructions, and instructions that work for larger models like GPT-4 may not suit smaller models. Similarly, SLMs face challenges in identifying optimal system instructions in various tasks for training that differ from LLMs’. Therefore, Orca 2 (Mitra et al., 2023) proposes a cautious reasoning SLM training process: (1) starting with diverse tasks, (2) guided by the performance of Orca 1 (Mukherjee et al., 2023) to select system instructions, (3) writing task-specific instructions, and (4) employing Prompt Erasing to replace detailed instructions with generic ones, encouraging models to help SLMs learn not just task solutions but also deeper reasoning abilities. Orca 2 introduces a new dataset with approximately 817K instances, training on Llama-2-7B or Llama-2-13B (Touvron et al., 2023b) with 32 NVIDIA A100 GPUs over about 70 hours, demonstrating efficient instruction tuning compared to training from scratch. Evaluation on the ARC-Challenge shows Orac-2-7B achieves 78.41 accuracy, outperforming Llama-2-Chat-70B (67.66) and close to ChatGPT (84.73), showing that instruction tuning in Orca 2 can rival models 5-10 times its size.
@@
Instruction Tuning Domain SLMs for QA. Beyond instruction tuning for generic SLMs, tuning domain-specific SLMs is also crucial, as they provide specialized assistance where generic SLMs may underperform. Instruction-tuning generic SLMs can derive domain SLMs. We summarize some representatives in several domains. (1) In the financial domain, Phogat et al. (2024) transfer financial QA abilities from generic teacher LLMs such as GPT-4 (Achiam et al., 2023) to specialized student SLMs, such as Phi-3-Mini (Abdin et al., 2024), using datasets including FinQA (Chen et al., 2021a), ConvFinQA (Chen et al., 2022a), and TATQA (Zhu et al., 2021). They apply step-wise arithmetic annotation via a program of thought (PoT) (Chen et al., 2023c) prompting, preparing few-shot examples that cover concept understanding, formula writing, extracting relevant entities, and performing calculations. Incorrect codes and formats in teacher annotations are removed before fine-tuning SLMs with LoRA (Hu et al., 2021). Tests on FinQA show Phi-Mini closely matches GPT-4’s accuracy (77.59 vs. 77.51 zero-shot and 78.46 few-shot), improving fine-grained financial QA abilities, especially in concept understanding and entity extraction. (2) In the medical field, Guo et al. (2023a) enhance student SLMs, including domain-specific BioGPT (1.6B)  (Luo et al., 2022) and general Llama 7B (Touvron et al., 2023a), by fine-tuning on enriched PubMedQA (Jin et al., 2019) data. This enhancement is achieved by generating new samples or rewriting existing ones using teacher LLMs, which include the highly knowledgeable GPT-4 and the relatively weaker ChatGPT. The best SLM, with under 1.6 billion parameters, achieves 75.4% accuracy, surpassing GPT-4’s 74.4% in few-shot settings on the PubmedQA test sets. It demonstrates that LLMs effectively refine and diversify question-answer pairs, leading to enhanced performance in a significantly smaller model after fine-tuning. Additionally, there are three interesting insights: (i) Direct tests on PubMedQA reveal that BioGPT (1.6B) performs comparably or worse than the larger Llama 7B (0.594 vs 0.63 accuracy, 0.495 vs 0.387 macro-F1). (ii) When fine-tuned on the same PubMedQA training data, BioGPT (1.6B) notably outperforms Llama 7B, scoring 0.498 to Llama 7B’s 0.463 macro-F1, highlighting that domain knowledge helps domain SLMs learn faster than general models. (iii) Stronger teacher LLMs improve fine-tuning outcomes. For example, BioGPT (1.6B) fine-tuned with new QA data from ChatGPT and GPT-4 shows that the latter setup, with a macro-F1 of 0.520, outperforms the former’s 0.498, indicating that more knowledgeable teachers significantly enhance SLM specialization. We report the detailed results of comparisons of instruction-tuned domain-specific language models (SLMs) for QA and larger language models on FinQA (Chen et al., 2021a) and PubMedQA (Jin et al., 2019), as shown in Table 5.
@@
Model Size Task Name Shot Type Accuracy (%)
GPT-4 (Achiam et al., 2023) - FinQA Zero-shot 77.5
Phi-3-Mini (Abdin et al., 2024) 2.7B FinQA Zero-shot 77.6
Meditron-70B (Chen et al., 2023a) 70B PubMedQA Zero-shot 81.6
RankRAG-llama3-70B (Yu et al., 2024b) 70B PubMedQA Zero-shot 79.8
Flan-PaLM (Singhal et al., 2023) 540B PubMedQA Few-shot 79.0
GAL 120B (Taylor et al., 2022) 120B PubMedQA Zero-shot 77.6
Flan-PaLM (Singhal et al., 2023) 62B PubMedQA Few-shot 77.2
BioGPT (Luo et al., 2022) 345M PubMedQA Zero-shot 78.2
BioGPT-Large (Luo et al., 2022) 1.5B PubMedQA Zero-shot 81.0
Enhancing SLMs for Out-of-Domain Questions. One of the major advantages of LLMs is their strong comprehension and logical reasoning abilities, which SLMs often struggle to match due to their limited parameters, especially when handling unseen or out-of-domain questions. Various methods have been developed to address this limitation, including Chain-of-Thought (CoT) prompting, Retrieval-Augmented Generation (RAG), and self-adaptive techniques. We categorize these techniques into the following groups:

### v2

@@ 3.2. Supervised Fine-Tuning (SFT) for Enhancing SLM performance
Supervised Fine-Tuning (SFT) employs a training methodology similar to pre-training but is specifically tailored to align models to adhere to the instructions encapsulated within various instructional datasets. This approach is designed to refine the model’s responsiveness and appropriateness to given contexts as dictated by the training data. For example, various models, such as Alpaca (Taori et al., 2023), UltraChat (Ding et al., 2023), WizardLM (Xu et al., 2023b), SlimOrca (Lian et al., 2023), ShareGPT (Wang et al., 2024d), Capybara (Daniele and Suphavadeeprasit, 2023), Deita (Liu et al., 2023c), and MetaMathQA (Yu et al., 2024a), incorporates a suite of conversational datasets to enhance their capabilities in context-aware dialogue and instruction adherence. Usually, as shown in Figure 11, existing SFT methods can be categorized into three categories:
  • (i) Classical fine-tuning with downstream data (Devlin et al., 2019; Radford et al., 2019) trains SLMs on task-specific annotated data, transferring general language representations to specific tasks such as sentiment analysis. In the LLM era, this approach remains effective, such as enhancing LLMs by calibrating responses or assigning risk scores with smaller models such as BERT (Zhao et al., 2023b), or optimizing for mobile devices with MobileBERT (Sun et al., 2020).
  • (ii) Instruction tuning with LLM-generated data (Taori et al., 2023; Ding et al., 2023; Lian et al., 2023) or human-generated questions with LLM annotations (Wang et al., 2024d) aims to align generative models with specific instructions, enhancing their instruction-following and reasoning capabilities. For example, Alpaca 7B (Taori et al., 2023) uses 52k ChatGPT-generated instruction-following examples from 175 self-instructed seed tasks to tune Llama 7B (Touvron et al., 2023a). Meanwhile, StableLM (Bellagente et al., 2024; Tow et al., [n. d.]) is trained on the Restruct-v1 dataset, which includes summarization, question-answering, and sentiment analysis tasks, using instruction data from (Longpre et al., 2023b).
  • (iii) Preference optimization with human feedback (Rafailov et al., 2024; Wang et al., 2024d; Ouyang et al., 2022) aims to better align language models with human preferences. Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) gathers human-preferred data, trains a reward model, and fine-tunes the LM using reinforcement learning. Direct Preference Optimization (DPO) (Rafailov et al., 2024) provides a simpler alternative to RLHF. Unlike RLHF, DPO avoids explicit reward modeling and reinforcement learning techniques. Instead, it adjusts the log probabilities of preferred versus non-preferred responses using a dynamic weighting mechanism, preventing model degradation issues typical of methods relying on probability ratios. For instance, Llama 3.2 1B & 3B apply SFT and DPO in post-training to enhance alignment with instructions and human preferences.
@@ 3.3. Data Quality in Knowledge Distillation (KD)
Augment Data from LLMs. LLM-generated data could be categorized as pre-training data and fine-tuning data. Firstly, due to the limitations of model size, studies have shown that training SLMs requires simple and comprehensible data (Eldan and Li, 2023; Xi et al., 2024; Lee et al., 2024b; Li et al., 2023c). As shown in Figure 12 (1) left, TinyStory (Eldan and Li, 2023) shows that small models (tens of millions of parameters) can generate coherent stories for 3-4-year-olds. GPT-3.5 or GPT-4 (Achiam et al., 2023) prompts create simple stories from three keywords chosen from a 1,500-word vocabulary, which are then used to train SLMs for similar outputs. This approach shows that simple and comprehensible data can help smaller models exhibit behaviors similar to those of larger language models, such as obeying scaling laws and achieving enhanced performance. On the other hand, many efforts to enhance the Chain-of-Thought (CoT) capabilities of small models involve using LLMs to generate high-quality fine-tuning CoT data. As shown in Figure 12 (1) right, these data train small models end-to-end to mimic CoT reasoning (Xi et al., 2024; Ma et al., 2023b). AS-ES Learning (Xi et al., 2024) highlights that small models struggle with complex reasoning, even when provided detailed steps, as these require nuanced extraction and abstraction. Therefore, the study introduces a paradigm splitting reasoning into extractive segments (context reminders) and abstractive segments (inferred insights).
@@
Figure 13. Distillation Techniques for Enhancing SLM Performance. On-policy means learning only use data from the current student (policy), while off-policy permits the use of previously gathered data.
@@ 3.5. Performance Improvement through Quantization
Quantization-Aware Training (QAT) differs from PTQ in that it includes a training phase after the model has been quantized. When models are quantized to extremes, such as 2-bit or 1-bit, performance typically drops significantly, but further training can help the model retain its capabilities. For instance, to mitigate performance degradation from binarization, PB-LLM (Shang et al., 2023) selectively binarizes only non-salient weights, preserving the most salient ones at higher precision. This method effectively reduces the model size without significantly impacting performance. Salient weights are chosen based on their magnitude, ensuring that the most influential weights maintain higher precision to preserve the model’s reasoning capabilities. The paper explores both post-training quantization (PTQ) and quantization-aware training (QAT) to fine-tune and recover the performance of partially binarized models, achieving a balance between compression and accuracy. OneBit (Xu et al., 2024b) and BitNet (Wang et al., 2023e) address the severe performance degradation associated with 1-bit quantization by decomposing floating-point matrices and employing mixed-precision strategies. Specifically, OneBit introduces Sign-Value-Independent Decomposition (SVID), which decomposes a floating-point matrix into a 1-bit matrix and two floating-point vectors. This method allows LLMs to be quantized to a 1-bit level while minimizing performance loss. By retaining critical information with the floating-point vectors, OneBit effectively balances extreme compression with maintaining model accuracy. BitNet b1.58 (Ma et al., 2024) improves on the original BitNet by introducing a ternary matrix weight system -1, 0, 1, resulting in a 1.58-bit model. BitNet b1.58 matches the performance of full-precision models starting from a 3 billion parameter size while further reducing memory and latency costs. LLM-QAT (Liu et al., 2023a) employs data-free knowledge distillation, where the pre-trained model itself generates data for fine-tuning the quantized model (student) using logit distillation from the full-precision model (teacher). This method incorporates quantization of weights, activations, and key-value cache, achieving accurate 4-bit quantization for weights and key-value caches, and 6-bit for activations, demonstrating substantial improvements over existing post-training quantization methods.
@@ 3.6. Techniques in LLMs Contributing to SLMs
This subsection explores the potential of advanced techniques such as RAG and MoE, which enhance LLM performance, to also maintain or boost SLM performance within constrained computational budgets. However, effectively integrating these techniques into SLMs, which inherently possess limited capabilities, remains an unresolved challenge.
Retrieval Augmented Generation (RAG) enhances the capabilities of language models in knowledge-intensive tasks by incorporating a retrieval mechanism. This approach allows models to access relevant contextual information from a data repository in response to user queries. By integrating this retrieved data, RAG-equipped models better understand specific topics, enabling more informed and accurate outputs. For SLMs, a significant concern is whether they possess the capacity for long-context reasoning. A recent study (Liu et al., 2024g) compares SLMs at the 7B level with RAG to larger models such as GPT-3.5 and GPT-4, suggesting that SLMs equipped with RAG can sometimes perform comparably or even better than LLMs. These findings indicate that RAG for SLMs is effective and represents a promising direction for future research.
Mixture-of-Experts (MoE) (Cai et al., 2024b) has emerged as an effective method for substantially scaling up model capacity with minimal computation overhead in LLMs. The MoE framework is founded on a straightforward yet potent concept: distinct components of a model, referred to as “experts”, specialize in different tasks or data facets. In this paradigm, only the relevant experts are activated for a specific input, which manages computational costs while leveraging a vast pool of specialized knowledge. This scalable and adaptable approach enables increased model capacity without proportionally escalating computational demands. We argue that MoE is particularly suitable for SLM architectures (Jiang et al., 2024a) as it minimizes both computational load and memory overhead. However, research on MoE for SLMs remains sparse. Future studies could investigate how large LLM MoE architectures can be effectively compressed into small ones or how to develop an SLM with MoE tailored for specific devices from scratch.
@@ 4. Applications of Small Language Models
In this section, we delve into the applications of small language models (SLMs) across various NLP tasks and their deployment strategies. Due to benefits such as enhanced privacy, faster inference, and lower memory requirements, many NLP applications are now leveraging SLMs over LLMs. Additionally, deploying SLMs often involves considerations of memory and runtime efficiency, which are crucial for optimizing resource use on budget-constrained edge devices, particularly mobile phones. Then, we will discuss task-specific applications of SLMs and their deployment methods on mobile and edge devices.
@@ 4.1. Task-specific SLM Applications
SLM in QA Alpaca (Taori et al., 2023) Tune Llama 7B (Touvron et al., 2023a) using 52k ChatGPT-generated examples.
Stable Beluga 7B (Mahan et al., [n. d.]) Employ explanation tuning to Llama-2 7B (Touvron et al., 2023b) on an Orca-style dataset.
Fine-tuned BioGPT (Guo et al., 2023a) Fine-tuning BioGPT (1.6B) (Luo et al., 2022) on PubMedQA.
Financial SLMs (Phogat et al., 2024) Transfer financial knowledge from GPT-4 (Achiam et al., 2023) to multiple SLMs.
ColBERT (Gichamba et al., 2024) Fetch retrieval documents for SLMs to answer domain-specific questions.
Rationale Ranking (Hartill et al., 2023a) For unseen questions, combine retrieval with LLM-generated rationales.
@@
SLM in Coding Phi-3.5-mini (Abdin et al., 2024) New addition to the Phi-3 series and focus on high-quality data.
TinyLlama (Zhang et al., 2024j) 1.1B Transformer model is trained on 3T corpus.
CodeLlama (Roziere et al., 2023) A derivative of Llama 2 fine-tuned on domain-specific datasets.
CodeGemma (Team, 2024a) Fine-tuning Gemma to enhancing coding capabilities.
@@
SLIM (Wang et al., 2024g) Step-by-step Knowledge Distillation
@@
SLM in Web Search Content encoder (Chang et al., [n. d.]; Humeau et al., [n. d.]; Lu et al., 2020) Encode concatenated queries and documents.
Ranker (Chu et al., 2022; Nogueira and Cho, 2019) Re-rank retrieved documents using a specially SLM.
Rewriter (Ma et al., 2023d) Bridge the gap between queries and needed knowledge by rewriting inputs.
@@
$\alpha$-UMI (Shen et al., 2024a) SLMs serve as Multi-agents in tool uses.
Mobile Interaction (Carreira et al., 2023) Text-to-action control and tests on 6GB and 4GB Android devices
@@
M4 (Yuan et al., 2024c) a foundation model handling all mobile AI tasks.
Agent for Text Rewriting (Zhu et al., 2023b) Data Knowledge Distillation from LLMs
@@ 4.1.1. SLM Applications in Question-Answering
Question-answering (QA) is a fundamental task in the NLP field, demanding language models to exhibit abilities in understanding language, reasoning, common sense, and recalling specialized knowledge. Typically, larger language models yield better QA performance. However, the substantial size of these models introduces challenges such as immense computational requirements, privacy concerns when using proprietary LLMs, and difficulties in customization. These issues lead researchers and developers to favor SLMs in scenarios that demand efficiency, privacy, and customization. Therefore, we explore methods to enhance the capabilities of SLMs in QA across three key areas: (i) Instruction Tuning of Generic SLMs for QA, (ii) Instruction Tuning of Domain-Specific SLMs for QA, and (iii) Enhancing SLMs for Out-of-Domain Questions.
@@
Instruction Tuning Generic SLMs for QA. Despite the Phi series’ high question-answering capability, its training cost with over 3.4T tokens on 512 H100 GPUs for 10 days (Abdin et al., 2024) is prohibitive for many researchers and developers. Instruction tuning (Wei et al., 2022) offers a cost-effective alternative, enhancing small models by fine-tuning on large model outputs. Alpaca 7B (Taori et al., 2023) tunes Llama 7B (Touvron et al., 2023a) with 52k ChatGPT-generated examples from 175 seed tasks. This behavior cloning mimics teacher models effectively but struggles in reasoning-intensive QA tasks where accuracy is key, not style (Chia et al., 2024). To counter it, explanation tuning (Mahan et al., [n. d.]) enhances Llama-2 7B (Touvron et al., 2023b) using explanatory LLM answers to improve reasoning. However, its effectiveness varies with system instructions, and those effective for larger models like GPT-4 may not suit smaller ones. SLMs also struggle to identify optimal system instructions for different tasks. Therefore, Orca 2 (Mitra et al., 2023) addresses this by promoting cautious reasoning, deciding which solution strategy to choose for a given task among direct answer generation, or “Slow Thinking” strategies (step-by-step, guess and check or explain-then-answer, etc.) and erasing specific system instructions during training. This involves (1) solution strategy is guided by the performance of Orca 1 (Mukherjee et al., 2023), (2) writing task-specific system instructions corresponding to the chosen strategy to obtain teacher responses for each task, and (3) at training time, employing Prompt Erasing to replace student’s system instructions with generic ones vacated of details of how to approach the task, encouraging students learn not just task solutions but also deeper reasoning abilities.
@@
Instruction Tuning Domain SLMs for QA. Beyond instruction tuning for generic SLMs, tuning domain-specific SLMs is also crucial, as they provide specialized assistance where generic SLMs may underperform. Instruction-tuning generic SLMs can derive domain SLMs. We summarize some representatives in several domains. (1) In finance, Phogat et al. (2024) transfer financial QA abilities from teacher LLMs such as GPT-4 (Achiam et al., 2023) to specialized SLMs such as Phi-3-Mini (Abdin et al., 2024), using datasets such as FinQA (Chen et al., 2021a), ConvFinQA (Chen et al., 2022a), and TATQA (Zhu et al., 2021). They train SLMs with Python programs created by the teacher model, which detail steps for financial reasoning, including concept comprehension, formula identification, entity extraction, and calculations. During inference, SLMs generate Python code that an external interpreter executes. (2) In the medical field, Guo et al. (2023a) enhance student SLMs, including domain-specific BioGPT (1.6B)  (Luo et al., 2022) and general Llama 7B (Touvron et al., 2023a), by fine-tuning on enriched PubMedQA (Jin et al., 2019) data. This enhancement is achieved by generating new samples or rewriting existing ones using teacher LLMs, which include the highly knowledgeable GPT-4 and the relatively weaker ChatGPT. The best SLM, with under 1.6 billion parameters, achieves 75.4% accuracy, surpassing GPT-4’s 74.4% in few-shot settings on the PubmedQA test sets. It demonstrates that LLMs effectively refine and diversify question-answer pairs, leading to enhanced performance in a significantly smaller model after fine-tuning. We report the detailed results of comparisons of instruction-tuned domain-specific language models for QA and larger language models on FinQA (Chen et al., 2021a) and PubMedQA (Jin et al., 2019), as shown in Table 4.
@@
Model Size Instruction tuned? Task Name Shot Type Accuracy (%)
GPT-4 (Achiam et al., 2023) - $\times$ FinQA Zero-shot 77.5
Phi-3-Mini (Abdin et al., 2024) 2.7B $\checkmark$ FinQA Zero-shot 77.6
Meditron-70B (Chen et al., 2023a) 70B $\times$ PubMedQA Zero-shot 81.6
RankRAG-llama3-70B (Yu et al., 2024b) 70B $\times$ PubMedQA Zero-shot 79.8
Flan-PaLM (Singhal et al., 2023) 540B $\times$ PubMedQA Few-shot 79.0
GAL 120B (Taylor et al., 2022) 120B $\times$ PubMedQA Zero-shot 77.6
Flan-PaLM (Singhal et al., 2023) 62B $\times$ PubMedQA Few-shot 77.2
BioGPT (Luo et al., 2022) 345M $\checkmark$ PubMedQA Zero-shot 78.2
BioGPT-Large (Luo et al., 2022) 1.5B $\checkmark$ PubMedQA Zero-shot 81.0
Enhancing SLMs for Out-of-Domain Questions. One of the major advantages of LLMs is their strong comprehension and logical reasoning abilities, which SLMs often struggle to match due to their limited parameters, especially when handling unseen or out-of-domain questions. Various methods have been developed to address this limitation, including Retrieval-Augmented Generation (RAG) and self-adaptive techniques.